# Constraints on Dark Matter with Future MeV Gamma-Ray Telescopes

### v2

@@ Abstract
A number of new balloon or space-based gamma-ray observatories have been proposed to close a “MeV gap” in sensitivity to gamma rays in the MeV-GeV energy band. One aspect of the science case for these instruments is their ability to constrain or discover decaying or annihilating dark matter. In this work, we forecast the sensitivity of these instruments for dark matter annihilation or decay for a range of possible Standard Model final states and compare to existing bounds.
@@ I Introduction
There are many different experimental approaches to detecting interactions between dark matter and the known particles of the Standard Model. One common distinction is between “direct detection”, where dark matter interacts directly with the apparatus, and “indirect detection”, where the apparatus detects signs of dark matter, such as its annihilation or decay products. Historically, direct detection efforts have been primarily focused on GeV-TeV range WIMPs. However, the past decade has seen an increase in well-motivated sub-GeV dark matter candidates, which has spurred an increase in interest in the direct detection of sub-GeV dark matter, resulting in many promising new detector technologies (e.g. Ref. [1] and references therein).
On the indirect detection front, however, there is still a notable “sensitivity gap” for gamma rays in the MeV range, which could be produced by the annihilation or decay of modestly heavier dark matter. To combat this, many instruments capable of detecting MeV-range gamma rays have been proposed. These include COSI, which has already been selected to run (for indirect-detection-sensitivity forecasts for COSI, see [2]), as well as GECCO, e-ASTROGAM, AMEGO, AdEPT, PANGU, GRAMS, and MAST. Ref. [3] explores the constraints that these instruments will be able to place on the annihilation cross-section of MeV-range dark matter under the Higgs and vector-portal models, while Ref. [4] explores the constraints that these instruments will be able to place on the fraction of dark matter composed of primordial black holes. Other works in the literature have explored the individual sensitivity of specific instruments to indirect-detection signals, e.g. [5, 6, 7].
In this work, we forecast model-independent constraints that these proposed instruments will be able to place on the dark-matter decay lifetime and annihilation cross-section for a range of Standard Model final states including photons, electrons, muons, and pions, and compare the forecasted sensitivity of these proposed instruments to existing constraints in the literature. Section II describes our calculations. The signal and background models that we assume are discussed in Sections II.1 and  II.2, respectively. In Section II.3, we describe the statistical method we used to derive our projected constraints from the signal and background models and the specifications of the instrument in question. In Section III, we describe each of the instruments considered in this work. We then present and discuss our results in Section IV; we also discuss systematic uncertainties and some internal consistency checks. We conclude in Section V. In Appendix A, we include some supplementary results for Draco and M31 targets. In Appendix B, we explore the effect of including atmospheric backgrounds.
@@ II Methods
In order to forecast constraints that these instruments may place on dark-matter decay and annihilation, we must determine what decay rates or annihilation cross-sections would generate a photon signal that is statistically distinguishable from background using these instruments as a function of the dark-matter mass. Since the parameters of our background models have some uncertainty, we use the Fisher-matrix method introduced in [8]. To use this method, we begin by describing models for the signal and background.
@@ II.1 Signal Modeling
The observed photon flux, as a function of photon energy, from the decay/annihilation of dark-matter particles of mass $m_{\chi}$ is given by
@@
where $a=1$ for decay and $a=2$ for annihilation. The factor in brackets is called the $D$-factor in the decay case or the $J$-factor in the annihilation case. The integral covers the entire observation region, which consists of all points in space that are within a particular angular distance to the center of the object from the point of view of the instrument. In this work, we consider an observation region of $10^{\circ}$ for both decay and annihilation. This is on the same order as the maximum angular resolution for the instruments under consideration (see Figure 2). Together, the first two factors represent the number of particles in the observation region available to decay, or the number of pairs of particles in the observation region available to annihilate. The third factor, $\Gamma$, represents the interaction rate per particle or pair of particles. This means that $\Gamma=1/\tau$ in the decay case and $\Gamma=\left<\sigma v\right>/(2f_{\chi})$ in the annihilation case, where $f_{\chi}=1$ if the particles are self-conjugate and $2$ otherwise. In this work, we assume the particles are self-conjugate. Finally, $dN/dE_{\gamma}(E_{\gamma})$ represents the distribution of photons per decay/annihilation event as a function of photon energy.
Table 2 lists the $J$- and $D$-factors used in this work. Table 1 lists the profile parameters used to derive these factors and the references from which these results are taken.
@@
Table 1: The NFW profile parameters used to compute our $J$- (annihilation) and $D$- (decay) factors, where $d$ is the distance from the Earth to the object and the other profile parameters are as in Equation 1.
@@
Table 2: $J$- (annihilation) and $D$- (decay) factors, computed using the NFW profile parameters in Table 1. The $J$-factors are in units of $\mathrm{MeV^{2}cm^{-5}}$, while the $D$-factors are in units of $\mathrm{MeVcm^{-2}}$.
Since this work focuses on MeV-range dark matter, we consider only the following six decay/annihilation final states: a photon pair, a photon and neutral pion, a neutral pion pair, an electron-positron pair, a muon/antimuon pair, and a pair of charged pions. In more realistic dark-matter models, interactions between dark matter and standard matter are more complicated (and can be analyzed using tools such as those developed in Refs. [12, 13]), but the final state will often be expressible as a linear combination of the states we consider in this work. Because neutral pions decay with a high branching ratio into $\gamma\gamma$, final states involving neutral pions can also stand in for situations where the annihilation/decay produces a new mediator that decays to two photons (although the results in this case will depend somewhat on the mediator mass if it is produced non-relativistically). In all six cases, we consider the final-state-radiation (FSR) photon spectrum as well as the photon spectra from any radiative decay channels with a branching ratio greater than 0.01. In the case of dark matter decaying or annihilating directly to two photons, we simply have
@@
In the case of dark matter decaying or annihilating to neutral pions, which decay directly to a photon pair, we simply have $dN/dE_{\pi}=2\delta(E_{\gamma}-m_{\pi}/2)$ in the rest frame of the neutral pions, where $E_{\pi}$ is the energy of the photon in the neutral pion’s frame. We now boost this signal to the observation frame, which we assume is stationary with respect to the original dark-matter particles. Let $E_{\chi}$ be the energy of the photon in the rest frame of the dark-matter particles. From Ref. [14] appendix B, we have (assuming isotropic scalar interactions)
@@
Plugging our delta-function signal into Equation 4, we obtain
@@
The FSR spectrum for dark-matter decay/annihilation into a lepton pair ($l\in\{e,\mu\}$) is given in [16] as follows:
@@
where $x\equiv 2E_{\gamma}/m_{\mu}$ and $r\equiv(m_{e}/m_{\mu})^{2}$. We use Equation 4, replacing $\pi$ with $\mu$, to boost this signal into the observation frame.
@@
Charged pions decay radiatively to muons (BR$>0.999$) or electrons (BR$<0.001$), with each of these processes producing photon spectra given in [16]. However, the dominant radiative decay signal is the signal from the subsequent muon decay into electrons given in Equation 15. This signal is boosted into the charged pion’s frame as follows:
@@
where $E_{\mu}$ is the photon’s energy in the muon’s frame, $E_{\pi}$ is the photon’s energy in the charged pion’s frame, $\epsilon^{\prime}\equiv m_{\mu}/m_{\pi}$, $t_{\mathrm{min}}=E_{\pi}\epsilon^{\prime}$, and $t_{\mathrm{max}}=\mathrm{min}(m_{\mu}/2,E_{\pi}/\epsilon^{\prime})$. After boosting into the charged pion’s frame, we then use Equation 4 to boost this signal into the dark matter’s rest frame.
We now have our photon flux as a function of dark-matter interaction rate for each of the decay/annihilation modes considered in this work. To model the data we can obtain from a real instrument with a nonzero energy resolution, we implement Gaussian smearing on a logarithmic scale and make the conservative assumption that the instruments have no chance of detecting photons immediately outside of their nominal energy range. This forms the function, with normalization parameterized by $\Gamma$, that we use as input to our Fisher matrix.
@@ II.2 Background Modeling
For all three of the targets we consider, we take the background models from [5]. For Draco and M31, we use the 2-parameter background model
@@
with fiducial parameters $A=2.4\times 10^{-3}\mathrm{\ MeV^{-1}cm^{-2}s^{-1}sr^{-1}}$, $\alpha=2$. This model captures the galactic and extragalactic backgrounds.
@@
For the Galactic Center, we model the galactic and extragalactic backgrounds separately. We split the galactic background into four components: a bremsstrahlung component, a $\pi^{0}$ component, an inverse-Compton-scattering (ICS) component fit to Fermi data at higher energies, and an ICS component fit to COMPTEL and EGRET data at lower energies. The spectrum of each of these components is given in Figure 2 in Ref. [6] for a $b\leq 5,l\leq 5$ ROI. In order to rescale each of these components to the $10^{\circ}$ cone ROI that we are considering, we use all-sky models of the same physical emission components to estimate the ratio of the integrated flux between the two ROIs. Specifically, for this purpose we use Model A from Ref. [17]. This model is broken up into two components: the ICS component and the $\pi^{0}$/bremsstrahlung component. For each of the two Model A components, we find the ratio of the flux from the $10^{\circ}$ cone ROI to the flux from the $b\leq 5,l\leq 5$ ROI, and this becomes our fiducial rescaling factor for that Model A component. These rescaling factors are close to independent of photon energy, so we simply compute them for 1 GeV photons. We then rescale the bremsstrahlung component and $\pi^{0}$ components from Ref. [6] by the $\pi^{0}$/bremsstrahlung rescaling factor from Model A, and we rescale the ICS components from Ref. [6] by the ICS rescaling factor from Model A. The rescaled plot is shown in Figure 1.
@@
We use an extragalactic background model of the same form as Equation 18, with fiducial parameters $A_{\text{e.g.}}=0.004135\mathrm{\ MeV^{-1}cm^{-2}s^{-1}sr^{-1}}$, $\alpha=2.8956$ [18]. This model underestimates the extragalactic background at energies above a few MeV (see Figure 10 of [19]), but is consistent with observation in the region where the extragalactic background is significant. In the fit, we vary the normalization $A_{\text{e.g.}}$ of this model but hold the slope $\alpha$ fixed.
@@
Figure 1: The background photon signal in the direction of the Galactic Center from Figure 2 in Ref. [6], rescaled to the $10^{\circ}$ cone ROI using data from Model A from Ref. [17] at 1 GeV.
@@
We only consider the irreducible backgrounds from cosmic gamma-rays in our analysis. In general, we also expect there to be instrumental backgrounds and backgrounds arising from the Earth’s atmosphere (especially for balloon experiments). However, most of the experiments we consider have not studied these backgrounds (and their ability to adjust sensitivity cuts to exclude them). For this reason, we defer such an analysis to a future point where more of the instrument concepts have been developed in depth (for example e-ASTROGAM has performed detailed background/instrument modeling in Ref. [20]), but caution that our results may be somewhat optimistic as a result. The GRAMS study in Ref. [21] finds atmospheric photon backgrounds somewhat larger than our Galactic Center background flux, but not by an enormous factor (within an order of magnitude or so).
@@ II.3 Statistical analysis
for the Galactic Center background model. Here each $A_{X}$ parameter describes a rescaling of the normalization for the component $X$ (as shown in Figure 1 for all the components except the extragalactic one), and the full background model is the sum of these components. The total differential flux is
@@
where $\vec{\theta}_{\mathrm{fid}}$ are the fiducial background parameters as given in the previous subsection with $\Gamma_{\mathrm{fid}}$ set to zero. $A_{\mathrm{eff}}$ represents the effective area of the instrument as a function of incident photon energy, and $T_{\mathrm{obs}}$ represents the projected observation time. In this work, we take $T_{\mathrm{obs}}=10^{6}\mathrm{\ s}$ for illustration (around 11.6 days or 280 hours, which could be achieved by a balloon flight with a large-field-of-view instrument or a pointed observation by a satellite). We note that the strength of our constraints scales with $\sqrt{T_{\mathrm{obs}}}$. We now have

### v3

@@ Abstract
A number of new balloon or space-based $\gamma$-ray observatories have been proposed to close a “MeV gap” in sensitivity to $\gamma$ rays in the MeV–GeV energy band. One aspect of the science case for these instruments is their ability to constrain or discover decaying or annihilating dark matter. In this work, we forecast the sensitivity of these instruments for dark matter annihilation or decay for a range of possible Standard Model final states and compare to existing bounds.
@@ I Introduction
There are many different experimental approaches to detecting interactions between dark matter and the known particles of the Standard Model. One common distinction is between “direct detection,” where dark matter interacts directly with the apparatus, and “indirect detection,” where the apparatus detects signs of dark matter, such as its annihilation or decay products. Historically, direct detection efforts have been primarily focused on GeV–TeV range weakly interacting massive particles. However, the past decade has seen an increase in well-motivated sub-GeV dark matter candidates, which has spurred an increase in interest in the direct detection of sub-GeV dark matter, resulting in many promising new detector technologies (e.g., Ref. [1] and references therein).
On the indirect detection front, however, there is still a notable “sensitivity gap” for $\gamma$ rays in the MeV range, which could be produced by the annihilation or decay of modestly heavier dark matter. To combat this, many instruments capable of detecting MeV-range $\gamma$ rays have been proposed. These include the Compton Spectrometer and Imager (COSI), which has already been selected to run (for indirect-detection-sensitivity forecasts for COSI, see [2]), as well as the Galactic Explorer with a Coded Aperture Mask Compton Telescope (GECCO), the enhanced ASTROGAM (e-ASTROGAM), the All-sky Medium Energy Gamma-ray Observatory (AMEGO), the Advanced Energetic Pair Telescope (AdEPT), the PAir-productioN Gamma-ray Unit (PANGU), the Gamma-Ray and AntiMatter Survey (GRAMS), and the Massive Argon Space Telescope (MAST). Reference [3] explores the constraints that these instruments will be able to place on the annihilation cross section of MeV-range dark matter under the Higgs and vector-portal models, while Ref. [4] explores the constraints that these instruments will be able to place on the fraction of dark matter composed of primordial black holes. Other works in the literature have explored the individual sensitivity of specific instruments to indirect detection signals, e.g. [5, 6, 7].
In this work, we forecast model-independent constraints that these proposed instruments will be able to place on the dark matter decay lifetime and annihilation cross section for a range of Standard Model final states including photons, electrons, muons, and pions and compare the forecasted sensitivity of these proposed instruments to existing constraints in the literature. Section II describes our calculations. The signal and background models that we assume are discussed in Secs. II.1 and  II.2, respectively. In Sec. II.3, we describe the statistical method we used to derive our projected constraints from the signal and background models and the specifications of the instrument in question. In Sec. III, we describe each of the instruments considered in this work. We then present and discuss our results in Sec. IV; we also discuss systematic uncertainties and some internal consistency checks. We conclude in Sec. V. In Appendix A, we include some supplementary results for Draco and M31 targets. In Appendix B, we explore the effect of including atmospheric backgrounds.
@@ II Methods
In order to forecast constraints that these instruments may place on dark matter decay and annihilation, we must determine what decay rates or annihilation cross sections would generate a photon signal that is statistically distinguishable from background using these instruments as a function of the dark matter mass. Since the parameters of our background models have some uncertainty, we use the Fisher-matrix method described in [8]. The Fisher information matrix characterizes the curvature of the likelihood near its maximum, where the likelihood is a function of the parameters governing the signal and background models (a review can be found in the context of cosmology in e.g., Ref. [9]). Inverting the Fisher matrix yields the covariance matrix for the relevant parameters, and so can be used to forecast the approximate statistical significance of a signal of a given normalization, if the likelihood is approximately Gaussian. This method quantifies the degeneracy between the dependence of our observed spectra on the signal and background parameters, allowing us to extract limits on the signal parameters that would be distinguishable from background uncertainty. To use this method, we begin by describing models for the signal and background.
@@ II.1 Signal modeling
The observed photon flux, as a function of photon energy, from the decay/annihilation of dark matter particles of mass $m_{\chi}$ is given by
@@
where $a=1$ for decay and $a=2$ for annihilation. The factor in brackets is called the $D$ factor in the decay case or the $J$ factor in the annihilation case. The integral covers the entire observation region, which consists of all points in space that are within a particular angular distance to the center of the object from the point of view of the instrument. In this work, we consider an observation region of $10^{\circ}$ for both decay and annihilation. This is on the same order as the maximum angular resolution for the instruments under consideration (see Fig. 2). Together, the first two factors represent the number of particles in the observation region available to decay or the number of pairs of particles in the observation region available to annihilate. The third factor, $\Gamma$, represents the interaction rate per particle or pair of particles. This means that $\Gamma=1/\tau$ in the decay case and $\Gamma=\left<\sigma v\right>/(2f_{\chi})$ in the annihilation case, where $f_{\chi}=1$ if the particles are self-conjugate and $2$ otherwise. In this work, we assume the particles are self-conjugate. Finally, $dN/dE_{\gamma}(E_{\gamma})$ represents the distribution of photons per decay/annihilation event as a function of photon energy.
Table 1 lists the profile parameters used to derive the $J$ and $D$ factors used in this work and the references from which these results are taken. Table 2 lists the $J$ and $D$ factors used in this work.
@@
Table 1: The NFW profile parameters used to compute our $J$ (annihilation) and $D$ (decay) factors, where $d$ is the distance from Earth to the object and the other profile parameters are as in Eq. (1).
@@
Table 2: $J$ (annihilation) and $D$ (decay) factors, computed using the NFW profile parameters in Table 1. The $J$ factors are in units of $\mathrm{MeV^{2}\ cm^{-5}}$, while the $D$ factors are in units of $\mathrm{MeV\ cm^{-2}}$.
Since this work focuses on MeV-range dark matter, we consider only the following six decay/annihilation final states: a photon pair, a photon and neutral pion, a neutral pion pair, an electron-positron pair, a muon/antimuon pair, and a pair of charged pions. In more realistic dark matter models, interactions between dark matter and standard matter are more complicated (and can be analyzed using tools such as those developed in Refs. [13, 14]), but the final state will often be expressible as a linear combination of the states we consider in this work. Because neutral pions decay with a high branching ratio into $\gamma\gamma$, final states involving neutral pions can also stand in for situations where the annihilation/decay produces a new mediator that decays to two photons (although the results in this case will depend somewhat on the mediator mass if it is produced nonrelativistically). In all six cases, we consider the final-state-radiation (FSR) photon spectrum as well as the photon spectra from any radiative decay channels with a branching ratio (BR) greater than 0.01. In the case of dark matter decaying or annihilating directly to two photons, we simply have
@@
In the case of dark matter decaying or annihilating to neutral pions, which decay directly to a photon pair, we simply have $dN/dE_{\pi}=2\delta(E_{\gamma}-m_{\pi}/2)$ in the rest frame of the neutral pions, where $E_{\pi}$ is the energy of the photon in the neutral pion’s frame. We now boost this signal to the observation frame, which we assume is stationary with respect to the original dark matter particles. Let $E_{\chi}$ be the energy of the photon in the rest frame of the dark matter particles. From Appendix B of Ref. [15], we have (assuming isotropic scalar interactions)
@@
Plugging our $\delta$-function signal into Eq. (4), we obtain
@@
The FSR spectrum for dark matter decay/annihilation into a lepton pair ($l\in\{e,\mu\}$) is given in [17] as follows:
@@
where $x\equiv 2E_{\gamma}/m_{\mu}$ and $r\equiv(m_{e}/m_{\mu})^{2}$. We use Eq. (4), replacing $\pi$ with $\mu$, to boost this signal into the observation frame.
@@
Charged pions decay radiatively to muons (BR$>0.999$) or electrons (BR$<0.001$), with each of these processes producing photon spectra given in [17]. However, the dominant radiative decay signal is the signal from the subsequent muon decay into electrons given in Eq. (15). This signal is boosted into the charged pion’s frame as follows:
@@
where $E_{\mu}$ is the photon’s energy in the muon’s frame, $E_{\pi}$ is the photon’s energy in the charged pion’s frame, $\epsilon^{\prime}\equiv m_{\mu}/m_{\pi}$, $t_{\mathrm{min}}=E_{\pi}\epsilon^{\prime}$, and $t_{\mathrm{max}}=\mathrm{min}(m_{\mu}/2,E_{\pi}/\epsilon^{\prime})$. After boosting into the charged pion’s frame, we then use Eq. (4) to boost this signal into the dark matter’s rest frame.
We now have our photon flux as a function of dark matter interaction rate for each of the decay/annihilation modes considered in this work. To model the data we can obtain from a real instrument with a nonzero energy resolution, we implement Gaussian smearing on a logarithmic scale and make the conservative assumption that the instruments have no chance of detecting photons immediately outside of their nominal energy range. This forms the function, with normalization parametrized by $\Gamma$, that we use as input to our Fisher matrix.
@@ II.2 Background modeling
For all three of the targets we consider, we take the background models from [5]. For Draco and M31, we use the two-parameter background model
@@
with fiducial parameters $A=2.4\times 10^{-3}\ \mathrm{\ MeV^{-1}\ cm^{-2}\ s^{-1}\ sr^{-1}}$, $\alpha=2$. This model captures the galactic and extragalactic backgrounds and has an estimated systematic error of 15% [16]. Since our analysis uses the square root of the background flux, this contributes an estimated systematic error in our analysis of about 7%.
@@
For the Galactic Center, we model the galactic and extragalactic backgrounds separately. We split the galactic background into four components: a bremsstrahlung component, a $\pi^{0}$ component, an inverse-Compton-scattering (ICS) component fit to Fermi data at higher energies, and an ICS component fit to Compton Telescope (COMPTEL) and the Energetic Gamma Ray Experiment Telescope (EGRET) data at lower energies. The spectrum of each of these components is given in Fig. 2 in Ref. [6] for a $b\leq 5^{\circ},l\leq 5^{\circ}$ region of interest (ROI). In order to rescale each of these components to the $10^{\circ}$ cone ROI that we are considering, we use all-sky models of the same physical emission components to estimate the ratio of the integrated flux between the two ROI. Specifically, for this purpose we use model A from Ref. [18]. This model is broken up into two components: the ICS component and the $\pi^{0}$/bremsstrahlung component. For each of the two model A components, we find the ratio of the flux from the $10^{\circ}$ cone ROI to the flux from the $b\leq 5^{\circ},l\leq 5^{\circ}$ ROI, and this becomes our fiducial rescaling factor for that model A component. These rescaling factors are close to independent of photon energy, so we simply compute them for 1 GeV photons. We then rescale the bremsstrahlung component and $\pi^{0}$ components from Ref. [6] by the $\pi^{0}$/bremsstrahlung rescaling factor from model A, and we rescale the ICS components from Ref. [6] by the ICS rescaling factor from model A. The rescaled plot is shown in Fig. 1.
@@
We use an extragalactic background model of the same form as Eq. (18), with fiducial parameters $A_{\text{e.g.}}=0.004135\ \mathrm{\ MeV^{-1}\ cm^{-2}\ s^{-1}\ sr^{-1}}$, $\alpha=2.8956$ [19]. This model underestimates the extragalactic background at energies above a few MeV (see Fig. 10 of [20]), but is consistent with observation in the region where the extragalactic background is significant. In the fit, we vary the normalization $A_{\text{e.g.}}$ of this model but hold the slope $\alpha$ fixed.
@@
Figure 1: The background photon signal in the direction of the Galactic Center from Fig. 2 in Ref. [6], rescaled to the $10^{\circ}$ cone ROI using data from model A from Ref. [18] at 1 GeV.
@@
We only consider the irreducible backgrounds from cosmic $\gamma$ rays in our analysis. In general, we also expect there to be instrumental backgrounds and backgrounds arising from Earth’s atmosphere (especially for balloon experiments). However, most of the experiments we consider have not studied these backgrounds (and their ability to adjust sensitivity cuts to exclude them). For this reason, we defer such an analysis to a future point where more of the instrument concepts have been developed in depth (for example e-ASTROGAM has performed detailed background/instrument modeling in Ref. [21]), but caution that our results may be somewhat optimistic as a result. The GRAMS study in Ref. [22] finds atmospheric photon backgrounds somewhat larger than our Galactic Center background flux, but not by an enormous factor (within an order of magnitude or so).
@@ II.3 Statistical analysis
for the Galactic Center background model. Here each $A_{X}$ parameter describes a rescaling of the normalization for the component $X$ (as shown in Fig. 1 for all the components except the extragalactic one), and the full background model is the sum of these components. The total differential flux is
@@
where $\vec{\theta}_{\mathrm{fid}}$ are the fiducial background parameters as given in the previous subsection with $\Gamma_{\mathrm{fid}}$ set to zero. $A_{\mathrm{eff}}$ represents the effective area of the instrument as a function of incident photon energy, and $T_{\mathrm{obs}}$ represents the projected observation time. In this work, we take $T_{\mathrm{obs}}=10^{6}\mathrm{\ s}$ for illustration (around 11.6 days or 280 h, which could be achieved by a balloon flight with a large-field-of-view instrument or a pointed observation by a satellite). We note that the strength of our constraints scales with $\sqrt{T_{\mathrm{obs}}}$. We now have